What is the nth term of the harmonic sequence 1/2,1/4,1/6,1/8
Answer by stanbon(75887)

What is the nth term of the harmonic sequence 1/2,1/4,1/6,1/8
-----
If n = 1, a(1) = 1/(2*1)
If n = 2, a(2) = 1/(2*2)
If n = 3, a(3) = 1/(2*3)
.....
If n = n, a(n) = 1/(2n)
----------------
Cheers,
Stan H.
-----------